# On the Moreau envelope properties of weakly convex functions

### v1

@@ 8 Second derivative of the Moreau envelope and convexity of the image of the proximal operator
For weakly convex functions, we now prove that the Moreau envelope is twice differentiable if $g$ is twice differentiable and Lipschitz on the image of the proximal operator. These second-order properties have been studied in previous works [30, 26, 25] in the convex setting. We will also show that the boundary of the convex envelope of $\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right)$ is of measure zero (Proposition 7). In other words, the image of the proximity operator of a weakly convex function is almost convex.
Lemma 15.
@@
By Proposition 4, $\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right)$ is an open set, so the Lipschitness and $\mathcal{C}^{2}$ properties of $g$ are well defined on this open set. Notice that Lemma 15 has been shown for convex functions in [29, Theorem 3.4].
Proof.
If $g$ is twice differentiable, by Proposition 4, we get that
@@
Combined with Proposition 5, we obtain
$\nabla g^{\gamma}(x)=\nabla g(\mathsf{Prox}_{\gamma g}\left(x\right)).$ (23)
@@
which shows the second part of Lemma 15. ∎
@@
Proof.
By Lemma 15, we have
@@
So $g^{\gamma}$ is $2L_{g}$-smooth on ${\mathbb{R}}^{d}$. ∎
@@
$\text{Leb}(\text{dom}(g)\setminus\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right))=0,$ (24)
@@
$\text{Leb}(\text{Conv}(\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right))\setminus\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right))=0,$ (25)
@@
Proposition 7 has been proved recently in [32, Lemma 1(iv)].
Proof.
We denote by $\text{dom}(g)$ the set where $g<+\infty$. By the weak convexity of $g$, we know that $\text{dom}(g)$ is convex. In fact, for $x,y\in\text{dom}(g)$ and $\lambda\in[0,1]$, $g+\frac{\rho}{2}\|\cdot\|^{2}$ is convex, so we have $g(\lambda x+(1-\lambda)y)\leq\lambda g(x)+\frac{\rho\lambda}{2}\|x\|^{2}+(1-\lambda)g(y)+\frac{\rho(1-\lambda)}{2}\|x\|^{2}-\frac{\rho}{2}\|\lambda x+(1-\lambda)y\|^{2}<+\infty$, and $\lambda x+(1-\lambda)y\in\text{dom}(g)$. By [23, Theorem 1], because $\text{dom}(g)$ is convex, we have that $\text{Leb}(\text{dom}(g)\setminus\text{int}(\text{dom}(g)))=0$.
By [15, Theorem 4.2.3], a convex function is differentiable almost everywhere on the interior of its domain. As $g+\frac{\rho}{2}\|\cdot\|^{2}$ is convex, it is thus differentiable almost everywhere on the interior of its domain. As a consequence $g$ is differentiable almost everywhere on the interior of its domain, $\text{int}(\text{dom}(g))$. Denoting as $\Gamma\subset\text{int}(\text{dom}(g))$ the subset on which $g$ is differentiable, we thus have $\text{Leb}(\text{int}(\text{dom}(g))\setminus\Gamma)=0$.
Moreover, for $x\in\Gamma$, $g$ is differentiable at $x$, so by Proposition 4, we have $x=\mathsf{Prox}_{\gamma g}\left(x+\gamma\nabla g(x)\right)\in\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right)$. So we get $\Gamma\subset\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right)$. Therefore, we can deduce that $\text{Leb}(\text{int}(\text{dom}(g))\setminus\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right))=0$.
@@
Moreover for $x\in{\mathbb{R}}^{d}$ and $y\in\text{dom}(g)$, by definition of the proximal operator, we get $g(\mathsf{Prox}_{\gamma g}\left(x\right))+\frac{1}{2\gamma}\|x-\mathsf{Prox}_{\gamma g}\left(x\right)\|^{2}\leq g(y)+\frac{1}{2\gamma}\|x-y\|^{2}$. So $g(\mathsf{Prox}_{\gamma g}\left(x\right))<+\infty$, which gives $\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right)\subset\text{dom}(g)$. Because $\text{dom}(g)$ is convex, we get that $\text{Conv}(\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right))\subset\text{dom}(g)$. Combining with relation (24), we obtain equation (25). ∎

### v2

@@ 8 Second derivative of the Moreau envelope and convexity of the image of the proximal operator
For weakly convex functions, we now prove that the Moreau envelope is twice differentiable if $g$ is twice differentiable and Lipschitz on the image of the proximal operator. These second-order properties have been studied in previous works [49, 40, 39] in the convex setting. We will also show that the boundary of the convex envelope of $\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right)$ is of measure zero (Proposition 8.7). In other words, the image of the proximity operator of a weakly convex function is almost convex.
Lemma 8.1.
@@
By Proposition 6.1, $\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right)$ is an open set, so the Lipschitness and $\mathcal{C}^{2}$ properties of $g$ are well defined on this open set. Notice that Lemma 8.1 has been shown for convex functions in [45, Theorem 3.4].
Proof 8.2.
If $g$ is twice differentiable, by Proposition 6.1, we get that
@@
Combined with Proposition 7.1, we obtain
(23) $\nabla g^{\gamma}(x)=\nabla g(\mathsf{Prox}_{\gamma g}\left(x\right)).$
@@
which shows the second part of Lemma 8.1.
@@
Proof 8.4.
By Lemma 8.1, we have
@@
So $g^{\gamma}$ is $2L_{g}$-smooth on ${\mathbb{R}}^{d}$.
Lemma 8.5.
Let $g$ be a $\rho$-weakly convex function and $\mathcal{C}^{k}$ in $x\in{\mathbb{R}}^{d}$ with $k\in{\mathbb{N}}^{\star}$. Then for $\rho\gamma<1$, $g^{\gamma}$ is $\mathcal{C}^{k}$ in $x\in{\mathbb{R}}^{d}$.
Lemma 8.5 is a generalization of [48, Theorem 3.12] in the weakly convex setting in.
Proof 8.6.
By Proposition 6.1 and Proposition 7.1, we have that
$\nabla g^{\gamma}=\frac{1}{\gamma}\left(I_{d}-\left(I_{d}+\gamma\nabla g\right)^{-1}\right).$
In the previous equation, the right-term is $\mathcal{C}^{k-1}$. Thus we obtain that $g^{\gamma}$ is $\mathcal{C}^{k}$.
@@
(24) $\text{Leb}(\text{dom}(g)\setminus\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right))=0,$
@@
(25) $\text{Leb}(\text{Conv}(\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right))\setminus\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right))=0,$
@@
Proposition 8.7 has been proved recently in [51, Lemma 1(iv)].
Proof 8.8.
We denote by $\text{dom}(g)$ the set where $g<+\infty$. By the weak convexity of $g$, we know that $\text{dom}(g)$ is convex. In fact, for $x,y\in\text{dom}(g)$ and $\lambda\in[0,1]$, $g+\frac{\rho}{2}\|\cdot\|^{2}$ is convex, so we have $g(\lambda x+(1-\lambda)y)\leq\lambda g(x)+\frac{\rho\lambda}{2}\|x\|^{2}+(1-\lambda)g(y)+\frac{\rho(1-\lambda)}{2}\|x\|^{2}-\frac{\rho}{2}\|\lambda x+(1-\lambda)y\|^{2}<+\infty$, and $\lambda x+(1-\lambda)y\in\text{dom}(g)$. By [37, Theorem 1], because $\text{dom}(g)$ is convex, we have that $\text{Leb}(\text{dom}(g)\setminus\text{int}(\text{dom}(g)))=0$.
By [27, Theorem 4.2.3], a convex function is differentiable almost everywhere on the interior of its domain. As $g+\frac{\rho}{2}\|\cdot\|^{2}$ is convex, it is thus differentiable almost everywhere on the interior of its domain. As a consequence $g$ is differentiable almost everywhere on the interior of its domain, $\text{int}(\text{dom}(g))$. Denoting as $\Gamma\subset\text{int}(\text{dom}(g))$ the subset on which $g$ is differentiable, we thus have $\text{Leb}(\text{int}(\text{dom}(g))\setminus\Gamma)=0$.
Moreover, for $x\in\Gamma$, $g$ is differentiable at $x$, so by Proposition 6.1, we have $x=\mathsf{Prox}_{\gamma g}\left(x+\gamma\nabla g(x)\right)\in\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right)$. So we get $\Gamma\subset\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right)$. Therefore, we can deduce that $\text{Leb}(\text{int}(\text{dom}(g))\setminus\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right))=0$.
@@
Moreover for $x\in{\mathbb{R}}^{d}$ and $y\in\text{dom}(g)$, by definition of the proximal operator, we get $g(\mathsf{Prox}_{\gamma g}\left(x\right))+\frac{1}{2\gamma}\|x-\mathsf{Prox}_{\gamma g}\left(x\right)\|^{2}\leq g(y)+\frac{1}{2\gamma}\|x-y\|^{2}$. So $g(\mathsf{Prox}_{\gamma g}\left(x\right))<+\infty$, which gives $\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right)\subset\text{dom}(g)$. Because $\text{dom}(g)$ is convex, we get that $\text{Conv}(\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right))\subset\text{dom}(g)$. Combined with (24), we obtain equation (25).
9 Numerical examples
In this section, we present various examples of weakly convex functions and their associated Moreau envelope.
Example 9.1.
We first study the function
$h(x)=\left\{\begin{array}[]{ll}\frac{1}{4}-\frac{1}{2}x^{2}&\mbox{if }|x|\leq\frac{1}{2}\\ \frac{1}{2}(|x|-1)^{2}&\mbox{otherwise.}\end{array}\right$.
$h$ is $\mathcal{C}^{1}$ and $1$-weakly convex and not convex. After some computation, we obtain that, for $\gamma\in(0,1)$
$h^{\gamma}(x)$ $=\left\{\begin{array}[]{ll}\frac{1}{4}-\frac{x^{2}}{2(1-\gamma)}&\mbox{if }|x|\leq\frac{1-\gamma}{2}\\ \frac{(|x|-1)^{2}}{2(1+\gamma)}&\mbox{otherwise.}\end{array}\right$.
$\mathsf{Prox}_{\gamma h}\left(x\right)$ $=\left\{\begin{array}[]{ll}\frac{x}{1-\gamma}&\mbox{if }|x|\leq\frac{1-\gamma}{2}\\ \frac{x+\gamma}{1+\gamma}&\mbox{if }x\geq\frac{1-\gamma}{2}\\ \frac{x-\gamma}{1+\gamma}&\mbox{otherwise.}\end{array}\right$.
On Figure 1, we present the graph of the function $h$ and $h^{\gamma}$ for some $\gamma\in(0,1)$, as well as the corresponding proximal operators. We observe that $h^{\gamma}$ is weakly convex but not convex. Moreover, the critical points and the global minimum are preserved.
Figure 1: Graph of the function $h$, $h^{\gamma}$ for $\gamma\in\{0.1,0.5,0.99\}$ and their associated proximal operators.
Example 9.2.
We study the function
$f(x)=\left\{\begin{array}[]{ll}\frac{1}{2}(x+1)^{2}&\mbox{if }x\leq-\frac{1}{2}\\ \frac{1}{4}-\frac{1}{2}x^{2}&\mbox{if }x\in[-\frac{1}{2},\frac{1}{4}]\\ \frac{3}{16}+\frac{1}{2}(x-\frac{1}{2})^{2}&\mbox{otherwise.}\end{array}\right$.
$f$ is $\mathcal{C}^{1}$ and $1$-weakly convex and not convex with an unique minimum global and three critical points. Its Moreau envelope writes, for $\gamma\in(0,1)$
$f^{\gamma}(x)$ $=\left\{\begin{array}[]{ll}\frac{(x+1)^{2}}{2(1+\gamma)}&\mbox{if }x\leq-\frac{1-\gamma}{2}\\ \frac{1}{4}-\frac{x^{2}}{2(1-\gamma)}&\mbox{if }x\in[-\frac{1-\gamma}{2},\frac{1-\gamma}{4}]\\ \frac{3}{16}+\frac{(x-\frac{1}{2})^{2}}{2(1+\gamma)}&\mbox{otherwise.}\end{array}\right$.
The corresponding proximal operator is
$\mathsf{Prox}_{\gamma f}\left(x\right)$ $=\left\{\begin{array}[]{ll}\frac{x-\gamma}{1+\gamma}&\mbox{if }x\leq-\frac{1-\gamma}{2}\\ \frac{x}{1-\gamma}&\mbox{if }x\in[-\frac{1-\gamma}{2},\frac{1-\gamma}{4}]\\ \frac{x+\frac{\gamma}{2}}{1+\gamma}&\mbox{otherwise.}\end{array}\right$.
On Figure 2, we present the graph of the function $f$ and $f^{\gamma}$ for some $\gamma\in(0,1)$. We observe that $f^{\gamma}$ is weakly convex but not convex. Moreover, the critical points and the global minimum are preserved.
Figure 2: Graph of the function $f$, $f^{\gamma}$ for $\gamma\in\{0.1,0.25,0.49\}$ and their associated proximal operators.
Example 9.3.
As a third example, we study the function
$u(x)=\frac{1}{2}\left(e^{-x}-x^{2}\right)$
$u$ is $\mathcal{C}^{\infty}$ and $1$-weakly convex and not convex. After some computation, we obtain that, for $\gamma\in(0,1)$
$u^{\gamma}(x)$ $=\frac{1-\gamma}{\gamma}\left(W\left(\frac{\gamma}{2(1-\gamma)}e^{-\frac{x}{1-\gamma}}\right)+\frac{1}{2}W^{2}\left(\frac{\gamma}{2(1-\gamma)}e^{-\frac{x}{1-\gamma}}\right)\right)-\frac{x^{2}}{2(1-\gamma)}$
$\mathsf{Prox}_{\gamma u}\left(x\right)$ $=\frac{x}{1-\gamma}+W\left(\frac{\gamma}{2(1-\gamma)}e^{-\frac{x}{1-\gamma}}\right),$
where $W$ is the Lambert function defined by $W(x)e^{W(x)}=x$.
In this case, $u^{\gamma}$ is $\mathcal{C}^{\infty}$, after computations we get
$\frac{\partial u^{\gamma}}{\partial x}(x)$ $=-\frac{1}{1-\gamma}\left(x+\frac{1-\gamma}{\gamma}W\left(\frac{\gamma}{2(1-\gamma)}e^{-\frac{x}{1-\gamma}}\right)\right)$
$\frac{\partial^{2}u^{\gamma}}{\partial^{2}x}(x)$ $=\frac{1}{1-\gamma}\left(\frac{2}{\gamma}\frac{W\left(\frac{\gamma}{2(1-\gamma)}e^{-\frac{x}{1-\gamma}}\right)}{1+W\left(\frac{\gamma}{2(1-\gamma)}e^{-\frac{x}{1-\gamma}}\right)}-1\right).$
We see that $\frac{\partial^{2}u^{\gamma}}{\partial^{2}x}\geq-\frac{1}{1-\gamma}$, so that $u^{\gamma}$ is $\frac{1}{1-\gamma}$-weakly convex, confirming the result of Lemma 5.7.
On Figure 3, we present the graph of the function $u$ and $u^{\gamma}$ for some $\gamma\in(0,1)$. We observe that $u^{\gamma}$ is weakly convex but not convex. Moreover, the critical points are preserved.
Figure 3: Graph of the function $u$, $u^{\gamma}$ for $\gamma\in\{0.1,0.5,0.8\}$ and their associated proximal operators.
Example 9.4.
We now introduce the function $\phi$ for $x\in{\mathbb{R}}$ by
(26) $\phi(x)=\left|x-2\lfloor\frac{x+1}{2}\rfloor\right|\in[0,1]$
and then study the function
$v(x)=\left\{\begin{array}[]{ll}\frac{1}{2}\phi(x)&\mbox{if }\phi(x)\in[0,\frac{1}{2}]\\ \frac{3}{8}-\frac{1}{2}(\phi(x)-1)^{2}&\mbox{otherwise.}\end{array}\right$.
The function $v$ is non differentiable on the set $\{2n|n\in{\mathbb{N}}\}$ and $1$-weakly convex and not convex. Moreover $v$ is $2$-periodic and even. By its definition in equation (6), it is clear that the Moreau envelope is also periodic and even. It writes, for $\gamma\in(0,1)$
$v^{\gamma}(x)$ $=\left\{\begin{array}[]{ll}\frac{\phi^{2}(x)}{2\gamma}&\mbox{if }\phi(x)\in[0,\frac{\gamma}{2}]\\ \frac{1}{2}\phi(x)-\frac{\gamma}{8}&\mbox{if }\phi(x)\in[\frac{\gamma}{2},\frac{1+\gamma}{2}]\\ \frac{3}{8}-\frac{(\phi(x)-1)^{2}}{2(1-\gamma)}&\mbox{otherwise}\end{array}\right$.
and its proximal operator is obtained as
$\mathsf{Prox}_{\gamma v}\left(x\right)$ $=\left\{\begin{array}[]{ll}0&\mbox{if }\phi(x)\in[0,\frac{\gamma}{2}]\\ \phi(x)-\frac{\gamma}{2}&\mbox{if }\phi(x)\in[\frac{\gamma}{2},\frac{1+\gamma}{2}]\\ \frac{\phi(x)-\gamma}{1-\gamma}&\mbox{otherwise.}\end{array}\right$.
On Figure 4, we present the graph of the function $v$ and $v^{\gamma}$ for some $\gamma\in(0,1)$. We observe that $v^{\gamma}$ is periodic, pair, weakly convex but not convex. Moreover, the critical points and global minimizers are preserved.
Figure 4: Graph of the function $v$, $v^{\gamma}$ for $\gamma\in\{0.1,0.5,0.99\}$ and associated proximal operators.
Example 9.5.
We now study the function $\psi=\frac{1}{2}\phi$ introduced in Equation (26) The function $\psi$ is non differentiable on the set $\{n|n\in{\mathbb{N}}\}$ and not weakly convex. Moreover $\psi$ is $2$-periodic and even. Therefore the Moreau envelope is also $2$-periodic and even. It writes, for $\gamma\in(0,1)$
$\psi^{\gamma}(x)$ $=\left\{\begin{array}[]{ll}\frac{\phi^{2}(x)}{2\gamma}&\mbox{if }\phi(x)\in[0,\frac{\gamma}{2}]\\ \frac{1}{2}\phi(x)-\frac{\gamma}{8}&\mbox{otherwise}\end{array}\right$.
and its proximal operator is obtained as
$\mathsf{Prox}_{\gamma\psi}\left(x\right)$ $=\left\{\begin{array}[]{ll}0&\mbox{if }\phi(x)\in[0,\frac{\gamma}{2}]\\ \phi(x)-\frac{\gamma}{2}&\mbox{otherwise.}\end{array}\right$.
On Figure 5, we present the graph of the function $\psi$ and $\psi^{\gamma}$ for some $\gamma\in(0,1)$. We observe that $\psi^{\gamma}$ is periodic, pair but not differentiable. Moreover, the critical points and global minimizers are preserved.
Figure 5: Graph of the function $\psi$, $\psi^{\gamma}$ for $\gamma\in\{0.1,0.5,0.99\}$ and associated proximal operators.
Example 9.6.
In many inverse problem [9, 7], the restoration task is reformulated as an optimization problem of the form
(27) $\inf_{x\in{\mathbb{R}}^{d}}\|\mathcal{A}(x)-y\|_{1}+\phi(x),$
with $\phi:{\mathbb{R}}^{d}\to{\mathbb{R}}$ a smooth regularization function, $\mathcal{A}:{\mathbb{R}}^{d}\to{\mathbb{R}}^{m}$ a degradation operator and $y\in{\mathbb{R}}^{m}$ an observation.
However, the previous objective function is not smooth. Therefore, the authors of [58, 8] propose to replace the convex $g=\|\cdot\|_{1}$ function by its Moreau envelope $g^{\gamma}=\frac{1}{\gamma}L_{\gamma}$, with $L_{\gamma}$ the Hubert loss [30] defined by
$L_{\gamma}(x)$ $=\left\{\begin{array}[]{ll}\frac{1}{2}\|x\|_{2}^{2}&\mbox{if }|x|<\gamma\\ \gamma(\|x\|_{1}-\frac{d\gamma}{2})&\mbox{otherwise,}\end{array}\right$.
where $d$ is the dimension of the space. It allows to run gradient descent algorithms to approximate solution of problem (27). The Moreau envelop has also been proposed in the same context for smoothing convex constraints [64]
Example 9.7.
A well-established line of research [53] use Langevin algorithm to sample a target distribution defined by
$\pi\propto e^{-f-g},$
with $f$ a data-fidelity term and $g$ a weakly-convex regularization. Such an algorithm can be written as
$x_{k+1}=x_{k}-\delta\nabla f(x_{k})-\delta\nabla g(x_{k})+\sqrt{2\delta}z_{k+1},$
with $z_{k+1}\sim\mathcal{N}(0,I_{d})$. However, in the case of non-differentiable regularization, this algorithm can not be run. Therefore, $g$ is approximated by $g^{\gamma}$ [46, 20, 41, 52, 14, 25], and thanks to Proposition 7.1, the previous algorithm can be written as
$x_{k+1}=x_{k}-\delta\nabla f(x_{k})-\frac{\delta}{\gamma}\left(x_{k}-\mathsf{Prox}_{\gamma g}\left(x_{k}\right)\right)+\sqrt{2\delta}z_{k+1}.$
This type of algorithm has recently shown state-of-the art performances for sampling image posterior distributions [52].
10 Conclusion
In this document, we have presented the main properties of the Moreau envelope for weakly convex function. We show that the Moreau envelope is regular in the sense that it is differentiable with respect to $\gamma\in(0,\frac{1}{\rho})$ and $x\in{\mathbb{R}}^{d}$, and that its derivatives can be explicitly computed. We highlight the intrinsic connections between the Moreau envelope and the convex conjugate. Moreover, we show that the Moreau envelope preserves key objects of interest in optimization, namely the minimizers and the critical points. Moreover, we study the second order properties and demonstrate that the image of the proximal operator is almost convex.
Acknowledgements
This study has been carried out with financial support from the French Direction Générale de l’Armement. We thanks Olivier Ley and Antonin Chambolle for the useful discussions and pointing us some references.